# Improving LLM Predictions via Inter-Layer Structural Encoders

### 24 Mar 2026 v1

@@ Abstract
The standard practice in Large Language Models (LLMs) is to base predictions on the final-layer token representations. Recent studies, however, show that intermediate layers encode substantial information, which may contain more task-relevant features than the final-layer representations alone. Importantly, it was shown that for different tasks, different layers may be optimal. In this work we introduce Inter-Layer Structural Encoders (ILSE), a powerful structural approach to learn one effective representation from the LLM’s internal layer representations all together. Central to ILSE is Cayley-Encoder, a mathematically grounded geometric encoder that leverages expander Cayley graphs for efficient inter-layer information propagation. We evaluate ILSE across 13 classification and semantic similarity tasks with 9 pre-trained LLMs ranging from 14 million to 8 billion parameters. ILSE consistently outperforms baselines and existing approaches, achieving up to 44% improvement in accuracy and 25% in similarity metrics. We further show that ILSE is data-efficient in few-shot regimes and can make small LLMs competitive with substantially larger models.
@@ 1 Introduction
Large Language Models (LLMs) have transformed natural language processing through their ability to learn rich, transferable representations (Vaswani et al., 2017; Devlin et al., 2019; Brown et al., 2020). It is common practice to use representations from the final model layer for task-specific prediction, like classification or similarity computation. This practice assumes that deeper layers contain the most task-relevant information, an assumption that recent works have challenged (Wallat et al., 2020; Fan et al., 2024; Skean et al., 2025).
Research into how information is represented across layers in language models revealed that different layers capture different linguistic features. Lower layers capture surface-level and part-of-speech information, middle layers encode syntactic structures, while upper layers focus on semantic relationships (Tenney et al., 2019; Jawahar et al., 2019; Clark et al., 2019). Wallat et al. (2020) and Skean et al. (2025) demonstrated that intermediate layer representations often capture substantial task-relevant information and can outperform final-layer representations, while the optimal layer remains task-dependent. A core question is therefore how to effectively combine representations from all layers.
Several methods have proposed aggregating representations from all layers to improve downstream performance. Peters et al. (2018) introduced ELMo, which learns a scalar weighted sum over all layer representations. While simple, this approach cannot capture complex interactions. More recently, ElNokrashy et al. (2024) proposed Depth-Wise Attention (DWAtt), which uses the final layer as a query to attend over previous layers. DWAtt enables non-linear combinations but anchors the attention to the final layer rather than treating all layers equally. Other approaches modify the transformer architecture itself (Pagliardini et al., 2024; Xiao et al., 2025), but require training the full architecture from scratch, which may be infeasible.
In this work, we propose Inter-Layer Structural Encoders (ILSE) - a rigorous, efficient and effective method for combining layer representations from all the layers of the LLM. For each layer, we compute a single representation by mean-pooling over tokens. ILSE then introduces a structure over these layer representations and uses DeepSets (Zaheer et al., 2017) or Graph Neural Networks (GNNs) (Gilmer et al., 2017) to learn how to combine them. A key consideration when using GNNs is the choice of graph structure. Recently, Bechler-Speicher et al. (2024) showed that GNNs tend to overfit the structure of the input graph, and therefore forming a graph with special structure where it does not necessarily carry information to the target variable, may result in overfitting and performance loss. Nonetheless, it was shown that regular graphs, i.e., graphs where all nodes have the same degree, are robust to this structural overfitting. Later, it was shown in Wilson et al. (2025) that a special family of regular graphs - Cayley graphs over the Special Linear Group $SL(2,\mathbb{Z}_{n})$, are also preferable in terms of information flow. Specifically, these graphs are bottleneck-free, meaning that all nodes can pass information between them efficiently (Alon and Yahav, 2021). Moreover, it was shown by Bechler-Speicher et al. (2025) that even when providing a molecular graph for molecular-property prediction, and it is likely that this structure carries information for the task, replacing it with a Cayley graph that is not tied to the task, can match or even boost performance. Inspired by these results, ILSE focuses on regular structures that are bottleneck-free, and specifically: (1) Cayley graphs over the Special Linear Group $SL(2,\mathbb{Z}_{n})$ (2) Sets, corresponding to empty graphs with no edges and (3) fully connected graphs.
@@
Over 13 classification and semantic text similarity (STS) tasks and 9 pre-trained LLMs ranging from 14 million to 8 billion parameters, we show that ILSE consistently outperforms existing approaches by large margins.
Our main contributions are:
@@
    We propose ILSE, a rigorous and mathematically grounded approach for learning a representation across all inter-layer representations of any LLM.
@@
    We demonstrate the effectiveness of ILSE in an extensive evaluation over 13 classification and STS tasks. Across all the classification tasks and the supervised STS task we evaluated, it outperforms all baselines, across all LLMs we evaluated.
@@
    In 7 zero-shot STS tasks we evaluated, ILSE outperforms all other baselines in 17 out of the total 21 combinations of LLM and tasks.
@@
    We demonstrate the data efficiency of ILSE in a few-shot evaluation, where with just 32 samples per label, it outperforms Last-Layer and Best-Layer baselines across all tasks. In addition, in one task using just 32 samples per label, ILSE outperforms all existing baselines trained on all the data (about 10,000 samples in total which are  130 samples per label).
@@
    We show that ILSE makes small LLMs on par with big LLMs, including making 14M-parameter LLM on par with a 2.8B-parameter LLM.
@@ 2 Related Work
2.1 Inter-Layer Representations in LLMs
@@ 3 Inter-Layer Structural Encoders
4 Experiments
  • Zero-shot STS transfer learning where we train on one STS task and evaluate the zero-shot performance across 7 other tasks.
  • Few-shot training with 1110241024 samples per label.
4.1 Tasks
Table 2: Performance comparison of ILSE and baselines across 3 different LLMs on 5 classification tasks. In all 15 cases ILSE gets the highest score. For Pythia-410m FC-Encoder is the most dominant method while in Gemma2-2B and Llama3-8B it’s Cayley-Encoder. In all cases but one ILSE also gets the best second and third scores. Bold: best per column, blue: 2nd best, red: 3rd best.
@@ 4.1 Tasks
4.2 Setup
  • MLP Last-Layer: A trained MLP over the last-layer representations.
  • MLP Best-Layer An MLP trained on the selected layer from Best-Layer.
Table 1 shows the number of learned parameters learned on top of the frozen LLM for each method. For the Last-Layer and Best-Layer methods, there are no added learned parameters.
Table 3: Performance comparison of ILSE and baselines across 3 different LLMs on 8 STS tasks. In 20 out of 24 tasks and LLMs, ILSE gets the best score, specifically Cayley-Encoder. Bold: best per column, blue: 2nd best, red: 3rd best.
@@ 4.2 Setup
For classification tasks, we train the encoders jointly with a linear classifier head using cross-entropy loss. The base LLM remains frozen throughout training. For STS tasks, we encode each sentence in a pair through the frozen LLM and our encoders, compute the cosine similarity between the resulting representations, and minimize Mean Squared Error (MSE) loss with respect to the ground-truth similarity score. We use the Adam optimizer (Kingma, 2014) with learning rate and weight decay selected via Optuna (Akiba et al., 2019) hyperparameter optimization on the validation set. For FC-Encoder and Cayley-Encoder, we evaluate two GNN architectures: GIN and GCN. We test different parameters values. For the number of MPNN layers we test 1-2, for dropout we test values 0.0 - 0.3, for learning rate and weight decay we test $10^{-4}$ to $10^{-3}$. For the number of layers in the MLP used inside the GIN aggregation we test 0-2. We keep hidden dimension fixed to 256, and batch size to 64 and 256 in classification and STS tasks respectively. We choose the best parameters set based on the best validation accuracy and train the final model. For the test evaluation we extract the trained encoder architecture without the classifier head for classification and use it for encoding as part of MTEB evaluation - we report the final accuracy and similarity scores on the test set.
@@
4.3 Results
Tables 2 and 3 present our main results. ILSE achieves the best performance in all 15 classification configurations, all 3 supervised STS configurations, and 17 out of 21 zero-shot STS transfer settings.
[…]
@@ 4.3 Results
4.4 Effectiveness Across LLM Scales
4.5 Few-Shot Learning
We evaluated the few-shot performance of ILSE by restricting the training data to a range of $1$ to $1024$ samples per label across all classification tasks (for some tasks where training data is limited, we reach the full training data after 128 samples per label).Our results indicate data efficiency. Over 4 tasks including Banking77, MTOP Domain, MTOP Intent and Poem Sentiment, ILSE achieves gains over both the last-layer and best-layer baselines with as few as $8$ samples per label. In Banking77 task using just $32$ samples per label makes Cayley-Encoder and FC-Encoder to surpass all other baselines including DWAtt being trained with full amount of training data. For Poem Sentiment task using 64 samples per label enables ILSE to outperform all other baselines being trained with more data. For Emotion task, starting with $32$ samples per label ILSE outperforms Last-Layer and Best-Layer baselines and for larger training set sizes ILSE outperforms all other baselines per training set size. Emotion is the only task that exhibits consistent performance improvements as the training set size grows for both ILSE and baselines. The other tasks reach plateau between $128$ to $512$ samples per label. The overall trend suggests that ILSE is effective in low-resource settings. This makes it particularly valuable for practical applications where labeled data is scarce (Figure 3).

### 28 Apr 2026 v2

@@ Abstract
The standard practice in Large Language Models (LLMs) is to base predictions on final-layer representations. However, intermediate layers encode complementary task-relevant signals, and the optimal layer is task-dependent, making single-layer usage inherently suboptimal. In this work, we introduce Inter-Layer Structural Encoders (ILSE), a powerful and parameter-efficient post-training framework that learns to aggregate representations from all layers of a frozen LLM through structured inter-layer interactions. Central to ILSE is the Cayley-Encoder, a mathematically grounded module based on expander Cayley graphs that enables efficient and effective inter-layer information propagation. We evaluate ILSE on 13 classification and semantic similarity tasks across 9 pre-trained LLMs ranging from 14M to 8B parameters. ILSE consistently outperforms strong baselines, achieving up to 44% improvements in accuracy and 25% in similarity, while introducing at most 0.1% additional parameters relative to the base LLM size. Furthermore, ILSE is highly data-efficient in few-shot regimes and enables small LLMs to match or exceed the performance of substantially larger models. Notably, it also outperforms LoRA-based fine-tuning despite operating on frozen representations.
@@ 1 Introduction
Large Language Models (LLMs) have transformed natural language processing through their ability to learn rich, transferable representations (Vaswani et al., 2017; Devlin et al., 2019; Brown et al., 2020). In practice, downstream predictions are typically based on the final-layer representations, under the assumption that deeper layers contain the most task-relevant information. However, recent work challenges this view, showing that intermediate layers encode complementary signals and can outperform the final layer depending on the task (Wallat et al., 2020; Fan et al., 2024; Skean et al., 2025).
A growing body of work has demonstrated that linguistic information is distributed across layers: lower layers capture surface-level features, middle layers encode syntactic structure, and higher layers focus on semantics (Tenney et al., 2019; Jawahar et al., 2019; Clark et al., 2019). Consequently, the optimal representation is inherently task-dependent, making reliance on a single layer suboptimal. This raises a fundamental question: how can representations from all layers be combined effectively? Several methods have proposed aggregating representations from all layers to improve downstream performance (Peters et al., 2018; ElNokrashy et al., 2024; Pagliardini et al., 2024; Xiao et al., 2025). However, these approaches either rely on heuristic or constrained aggregation mechanisms, or require modifying the transformer architecture and retraining it from scratch, incurring substantial computational cost.
In this work, we introduce Inter-Layer Structural Encoders (ILSE), a parameter-efficient post-training framework that aggregates representations from all layers of a frozen LLM into a single, improved representation. ILSE treats layer representations as elements of a structured object and learns their interactions using either DeepSets (Zaheer et al., 2017) or Graph Neural Networks (GNNs) (Gilmer et al., 2017).
In contrast to prior approaches, ILSE adopts a principled and mathematically grounded formulation rooted in recent advances in graph learning. In particular, it avoids overfitting to the chosen structure by leveraging regular graphs (Bechler-Speicher et al., 2024), while enabling efficient and bottleneck-free inter-layer communication via expander Cayley graphs (Alon and Yahav, 2021; Wilson et al., 2025). Concretely, ILSE employs structured inter-layer aggregation based on three regimes: (1) Cayley graphs over the Special Linear Group $SL(2,\mathbb{Z}_{n})$, which combine sparsity with strong connectivity guarantees, (2) fully connected graphs, and (3) sets.
@@
We evaluate ILSE on 13 classification and semantic text similarity (STS) tasks across 9 pre-trained LLMs ranging from 14M to 8B parameters. ILSE consistently outperforms existing approaches by large margins, while remaining highly parameter-efficient and data-efficient.
Our main contributions are as follows.
@@
    We propose ILSE, a principled and mathematically grounded, parameter-efficient post-training framework for structured aggregation of representations across all layers of a frozen LLM.
@@
    In an extensive evaluation of ILSE across 9 pre-trained LLMs and 13 classification and semantic similarity tasks, we show consistent improvements over strong baselines, achieving up to 44% improvements in accuracy and 25% in similarity, while introducing at most 0.1% additional parameters relative to the base LLM size.
@@
    We perform a few-shot evaluation with increasing amount of samples, we establish the data efficiency of ILSE, achieving strong performance with as few as 32 samples per label and outperforming full-data baselines in low-resource regimes.
@@
    We perform an evaluation of ILSE over a wide range of LLM sizes, ranging from 14 millionth 2.8 billion parameters, showing that it enables small LLMs to match or exceed the performance of substantially larger ones.
@@
    We show that ILSE outperforms LoRA-based fine-tuning.
@@ 2 Related Work
2.1 Aggregated Representations in LLMs
@@ 3 Inter-Layer Structural Encoders
4 Experiments
  • An evaluation of zero-shot transfer capabilities on STS tasks.
  • Few-shot training with 1110241024 samples per label.
  • A comparison with LoRA fine-tuning.
  • An evaluation of Cayley-Encoders on token-level instead of layer-level.
4.1 Tasks
4.2 Setup
  • MLP Last-Layer: A trained MLP over the last-layer representations.
  • MLP Best-Layer An MLP trained on the selected layer from Best-Layer.
Table 1 shows the number of learned parameters learned on top of the frozen LLM for each method. For the Last-Layer and Best-Layer methods, there are no added learned parameters. Our code and experimental configurations are publicly available. [^1] Code: https://github.com/tomulanovski/ILSE
For classification tasks, we train the encoders jointly with a linear classifier head using cross-entropy loss. The base LLM remains frozen throughout training. For STS tasks, we encode each sentence in a pair through the frozen LLM and our encoders, compute the cosine similarity between the resulting representations, and minimize Mean Squared Error (MSE) loss with respect to the ground-truth similarity score. We use the Adam optimizer (Kingma, 2014) with learning rate and weight decay selected via Optuna (Akiba et al., 2019) hyperparameter optimization on the validation set. For FC-Encoder and Cayley-Encoder, we evaluate two GNN architectures: GIN and GCN. We test different parameters values. For the number of MPNN layers we test 1-2, for dropout we test values 0.0 - 0.3, for learning rate and weight decay we test $10^{-4}$ to $10^{-3}$. For the number of layers in the MLP used inside the GIN aggregation we test 0-2. We keep hidden dimension fixed to 256, and batch size to 64. We choose the best parameters set based on the best validation accuracy and train the final model. For the test evaluation we extract the trained encoder architecture without the classifier head for classification and use it for encoding as part of MTEB evaluation - we report the final accuracy and similarity scores on the test set.
@@ 4.2 Setup
4.3 Results
Tables 2 and 4 present our main results. ILSE achieves the best performance in all 15 classification configurations, all 3 supervised STS configurations, and 17 out of 21 zero-shot STS transfer settings.
[…]
Table 2: Performance comparison of ILSE and baselines across 3 different LLMs on 5 classification tasks. For each task and LLM, we highlight in Bold the best performing setting, in blue the 2nd best and in red the 3rd best. In all 15 cases ILSE gets the highest score. For Pythia-410m FC-Encoder is the most dominant method while in Gemma2-2B and Llama3-8B it’s Cayley-Encoder. In all cases but one ILSE also achieves the best second and third scores.
@@ 4.3 Results
Since Cayley-Encoder maps layer representations randomly to nodes of the Cayley graph, we verify that performance is not sensitive to the specific permutation. We ran 10 independent random layer-to-node mappings with fixed model seeds on all classification tasks across all three LLMs. Table 3 reports mean and standard deviation across the 10 mappings, showing that performance remains stable with minimal variance, confirming that Cayley-Encoder’s gains are robust to the choice of layer-to-node assignment.
Table 3: Shuffled layer permutation results on Cayley graph topology. We report the mean $\pm$ std over 10 random layer permutations.
Base Model Method Banking77 Emotion MTOP Domain MTOP Intent Poem Sentiment
Pythia-410m Cayley-Encoder (GIN) $89.55\pm 0.34$ $74.91\pm 0.78$ $98.67\pm 0.06$ $94.55\pm 0.39$ $71.26\pm 1.83$
Cayley-Encoder (GCN) $90.07\pm 0.31$ $69.76\pm 1.53$ $98.57\pm 0.09$ $94.31\pm 0.40$ $68.05\pm 0.81$
Gemma2-2B Cayley-Encoder (GIN) $92.35\pm 0.19$ $71.50\pm 2.98$ $99.00\pm 0.08$ $96.22\pm 0.11$ $77.32\pm 2.89$
Cayley-Encoder (GCN) $92.32\pm 0.27$ $73.14\pm 1.46$ $99.14\pm 0.05$ $96.08\pm 0.17$ $82.25\pm 0.54$
Llama3-8B Cayley-Encoder (GIN) $92.35\pm 0.12$ $71.16\pm 0.71$ $98.88\pm 0.08$ $96.42\pm 0.08$ $74.46\pm 1.71$
Cayley-Encoder (GCN) $92.92\pm 0.13$ $72.80\pm 1.30$ $98.90\pm 0.08$ $96.08\pm 0.27$ $79.24\pm 1.31$
Table 4: Performance comparison of ILSE and baselines across 3 different LLMs on 8 STS tasks. For each task and LLM, we highlight in Bold the best performing setting, in blue the 2nd best and in red the 3rd best. In 20 out of 24 tasks and LLMs, ILSE gets the best score, specifically Cayley-Encoder.
@@
4.4 Effectiveness Across LLM Scales
@@
4.5 Few-Shot Learning
We evaluate the few-shot performance of ILSE by restricting the training data to a range of $1$ to $1024$ samples per label across all classification tasks (for some tasks where training data is limited, we reach the full training data after 128 samples per label).Our results indicate data efficiency. Over 4 tasks including Banking77, MTOP Domain, MTOP Intent and Poem Sentiment, ILSE achieves gains over both the last-layer and best-layer baselines with as few as $8$ samples per label. In Banking77 task using just $32$ samples per label makes Cayley-Encoder and FC-Encoder to surpass all other baselines including DWAtt being trained with full amount of training data. For Poem Sentiment task using 64 samples per label enables ILSE to outperform all other baselines being trained with more data. For Emotion task, starting with $32$ samples per label ILSE outperforms Last-Layer and Best-Layer baselines and for larger training set sizes ILSE outperforms all other baselines per training set size. Emotion is the only task that exhibits consistent performance improvements as the training set size grows for both ILSE and baselines. The other tasks reach plateau between $128$ to $512$ samples per label. The overall trend suggests that ILSE is effective in low-resource settings. This makes it particularly valuable for practical applications where labeled data is scarce (Figure 3).
@@ 4.5 Few-Shot Learning
4.6 Comparison with LoRA
Although ILSE is designed as a lightweight post-hoc module that is complementary to fine-tuning approaches and can be combined with methods such as LoRA (Hu et al., 2022) or full fine-tuning, we compare it directly against LoRA to demonstrate its standalone effectiveness. LoRA inserts trainable low-rank matrices into the LLM’s attention layers, adapting the model’s internal computation while keeping the original weights frozen. To keep the comparison fair in number of parameters, we capped the LoRa rank to 3. Tables 5 and 6 present the results on classification and STS tasks, respectively. On classification, ILSE outperforms LoRA in 13 out of 15 configurations of task and LLM combination. On STS tasks, Cayley-Encoder outperforms LoRA in 7 out of 8 tasks across all three LLMs. These results show the strength of structured inter-layer representations aggregation.
Table 5: ILSE vs LoRA accuracy on classification tasks. For each encoder, we report the best result between GIN and GCN variants. For each task and LLM, we highlight in Bold the best performing setting.
Base Model Method Banking77 Emotion MTOPDomain MTOPIntent PoemSentiment
Pythia-410m Set-Encoder 84.23 47.89 97.59 92.21 73.37
FC-Encoder 90.65 75.61 98.68 95.04 75.77
Cayley-Encoder 89.43 73.83 98.77 94.72 70.87
LoRA 79.35 91.06 96.64 88.25 70.96
Gemma2-2B Set-Encoder 90.03 78.77 98.92 94.37 78.56
FC-Encoder 92.39 79.90 99.07 95.34 77.12
Cayley-Encoder 92.58 69.60 99.16 96.43 83.27
LoRA 56.77 70.89 89.65 79.64 61.92
Llama3-8B Set-Encoder 87.62 71.04 98.77 95.43 77.02
FC-Encoder 92.38 71.64 98.99 96.19 77.98
Cayley-Encoder 92.85 73.43 99.03 96.46 79.04
LoRA 90.82 91.98 98.59 94.44 76.15
Table 6: ILSE vs LoRA spearman correlation ($\times 100$) on STS tasks. For each encoder, we report the best result between GIN and GCN variants of STSBenchmark. For each task and LLM, we highlight in Bold the best performing setting.
Base Model Method STSBenchmark STS12 STS13 STS14 STS15 STS16 BIOSSES SICK-R
Pythia-410m Set-Encoder $41.0$ $\mathbf{56.0}$ $43.7$ $44.2$ $44.0$ $36.0$ $47.0$ $42.6$
FC-Encoder $54.2$ $50.1$ $53.0$ $51.8$ $63.9$ $57.1$ $58.6$ $\mathbf{57.1}$
Cayley-Encoder $\mathbf{55.8}$ $56.0$ $\mathbf{60.2}$ $\mathbf{56.7}$ $\mathbf{67.0}$ $\mathbf{58.6}$ $56.6$ $55.3$
LoRA $54.1$ $44.6$ $51.7$ $52.3$ $64.8$ $45.8$ $\mathbf{64.3}$ $52.0$
Gemma2-2B Set-Encoder $36.7$ $44.0$ $39.0$ $35.1$ $50.5$ $47.3$ $31.6$ $42.2$
FC-Encoder $\mathbf{62.9}$ $\mathbf{61.8}$ $55.4$ $63.9$ $\mathbf{74.7}$ $\mathbf{63.0}$ $66.0$ $60.1$
Cayley-Encoder $61.6$ $60.4$ $\mathbf{64.0}$ $\mathbf{64.6}$ $73.9$ $62.7$ $64.1$ $\mathbf{60.3}$
LoRA $53.6$ $44.6$ $49.1$ $50.8$ $63.5$ $46.7$ $\mathbf{66.4}$ $52.2$
Llama3-8B Set-Encoder $42.5$ $42.0$ $50.6$ $49.1$ $50.0$ $41.2$ $34.3$ $44.7$
FC-Encoder $59.3$ $55.2$ $53.5$ $59.9$ $73.7$ $57.2$ $62.8$ $57.5$
Cayley-Encoder $\mathbf{63.1}$ $\mathbf{65.3}$ $\mathbf{63.5}$ $\mathbf{70.2}$ $\mathbf{77.0}$ $\mathbf{63.1}$ $55.3$ $\mathbf{60.7}$
LoRA $62.8$ $47.8$ $59.0$ $61.0$ $70.6$ $52.1$ $\mathbf{64.1}$ $59.1$
4.7 Token-Level Inter-Layer Aggregation
In all experiments above, each layer is represented by a single vector obtained via mean-pooling over tokens. While this is effective, it discards fine-grained token-level information. To investigate it, we construct a graph where each node corresponds to a token at a specific layer, resulting in $L\times T$ nodes per input, where $L$ is the number of layers of the base LLM and $T$ is the number of tokens in the input sequence. We apply Cayley-Encoder and FC-Encoder over this graph using Pythia-14M on three classification tasks . Table  7 shows substantial improvements over the mean-pooled variant. On Banking77, accuracy increases from 70–77% to 81–87%, on Emotion from 40–47% to 73–87%, and on Poem Sentiment from 48–55% to 50–58%. However, this comes at a significant computational cost. The number of graph nodes grows from $L$ to $L\times T$, substantially increasing memory and runtime requirements. This setting also highlights the scalability advantage of Cayley-Encoder over FC-Encoder. With $L\times T$ nodes, the fully-connected graph becomes large, with peak memory usage of 828–1516 MB compared to only 77–121 MB for Cayley-Encoder. This demonstrates that the sparse $O(n)$ edge complexity of the Cayley topology is highly beneficial when scaling to larger graphs in compare to the $O(n^{2})$ cost of FC-Encoder.
Table 7: Comparison of pooled-layer vs. all-tokens ILSE methods on Pythia-14m (L=7L{=}7 layers, D=128D{=}128). The original ILSE aggregated tokens via mean pooling per layer resulting, $L$ nodes; All-Tokens use every (layer, token) pair as a node ($L\times T$ nodes). All-tokens results show mean $\pm$ std over 3 seeds. Node ranges are estimated from 10 random samples per task. Peak memory is averaged across seeds and tasks. Cayley-Encoder shows lower peak memory - about 10 fold change.
Base Model Aggregation Method Banking77 Emotion PoemSentiment Nodes Peak Mem (MB)
Pythia-14m Mean-Pooled ($L$ nodes) Cayley-Encoder (GCN) $69.97$ $39.78$ $47.69$ 77
FC-Encoder (GCN) $77.12$ $47.46$ $54.81$ 77
Cayley-Encoder (GIN) $71.00$ $45.97$ $48.27$ 77
FC-Encoder (GIN) $75.54$ $47.30$ $36.73$ 77
All-Tokens ($L\times T$ nodes) Cayley-Encoder (GCN) $85.41\pm 0.47$ $87.88\pm 0.95$ $55.03\pm 0.06$ 2828175175 $121$
FC-Encoder (GCN) $81.74\pm 0.57$ $73.84\pm 1.38$ $50.32\pm 9.47$ 2828175175 1,5161{,}516
Cayley-Encoder (GIN) $7.46\pm 0.42 88$$8.26\pm 0.11$ $58.01\pm 4.69$ 2828175175 $77$
FC-Encoder (GIN) $81.21\pm 0.76$ $74.41\pm 0.93$ $58.17\pm 3.44$ 2828175175 $828$